OpenStudy (anonymous):

KMnO4 was made by dissolving 118.5g with water in a 250 mL volumetric flask. What is the molar concentration of the resulting solution?

OpenStudy (anonymous):

I'm nor sure how to set this up :/

OpenStudy (jfraser):

how many moles of KMnO4 is that 118.5g?

OpenStudy (callisto):

no. of mole of KMnO4 = 118.5/(39.1+54.9+16.0X4) = 0.75mole ; molar concentration = 0.75 / (250/1000) = 3moldm^-3 = 3M

OpenStudy (callisto):

JFraser, check the periodic table and find the molar mass of KMnO4 to calculate no. of mole of KMnO4!

OpenStudy (jfraser):

I'm well aware of how to do that, i was asking the OP because I wanted them to find it, but thanks anyway!